# Monitoring-feedback induced entanglement relaxations in a tilted free fermionic chain

### v1

@@ A.1 Numerical simulation for evolution of the state
In this section, we use the site index $m$ to replace the index $l$ which does not take into consideration the $L$-th site. Now we consider the continuous measurement with feedback and numerical simulation the quantum jump, i.e., $\left|\psi\left(t\right)\right\rangle\longrightarrow{\hat{L}_{m}\left|\psi\left(t\right)\right\rangle}/{\sqrt{\langle\hat{L}_{m}^{\dagger}\hat{L}_{m}\rangle}}$. The quantum jump for probability can be written as $p_{m}=\langle\hat{L}_{m}^{\dagger}\hat{L}_{m}\rangle\gamma\delta t$. Using Wick theorem, i.e., $\langle\hat{d}_{m}^{\dagger}\hat{d}_{m}\hat{d}_{m}^{\dagger}\hat{d}_{m}\rangle=\langle\hat{d}_{m}^{\dagger}\hat{d}_{m}\rangle\langle\hat{d}_{m}^{\dagger}\hat{d}_{m}\rangle+\langle\hat{d}_{m}^{\dagger}\hat{d}_{m}\rangle\langle\hat{d}_{m}\hat{d}_{m}^{\dagger}\rangle$, and fermionic anticommutation relations, i.e., $\{\hat{c}_{i},\hat{c}_{j}^{\dagger}\}=\delta_{ij}$, we get the probability of quantum jump, $p_{m}=\langle\hat{d}_{m}^{\dagger}\hat{d}_{m}\rangle\gamma\delta t=\left\langle\psi(t)\right|\hat{d}_{m}^{\dagger}\hat{d}_{m}\left|\psi(t)\right\rangle\gamma\delta t=|\hat{d}_{m}\left|\psi(t)\right\rangle|^{2}\gamma\delta t$. Note that the linear expansion of $\hat{d}_{m}$ can be written as $\hat{d}_{m}={\textstyle\sum_{l}^{L}{[d_{m}]_{l}\hat{c}_{l}}}$, where $d_{m}$ is a row vector. To simulate the quantum jump, we can find a single-particle state which satisfies $\hat{d}_{m}\left|\psi_{n}\right\rangle\neq 0$ and $n>1$, then move the corresponding column $U_{n}$ to the first due to gauge freedom and transform the rest columns of $U(t)$ by
$U_{n}^{\prime}=U_{n}-\frac{d_{m}U_{n}}{d_{m}U_{1}}U_{1}.$ (8)
@@
$U^{\prime}\left(t\right)=\begin{bmatrix}d_{m}^{\dagger}&U_{2}^{\prime}&U_{3}^{\prime}&\cdots&U_{N}^{\prime}\end{bmatrix}.$ (9)
Finally, we use the QR decomposition for $e^{i\pi\mathcal{N}_{m+1}}U^{\prime}(t)$, and $\mathcal{N}_{m+1}$ is the matrix form of particle number operator. It should be emphasized that, in our numerical simulation, $\delta t=0.01$ , $l=1,2,3,...,L$, $m=1,2,3,...,L-1$, $n=1,2,3,...,N$, except where other noted. Only with a sufficiently small time step can we better simulate the impact of continuous measurements on the evolution of the state.
A.2 Trajectory averaged observables
@@
When calculating the bipartite entanglement entropy $S_{L/2}$, we choose the sub system $[1,L/2]$ and then get the all eigenvalues $\lambda_{i}$} of sub-correlation matrix $C_{ij}^{L/2}$. $S_{L/2}$ is given by  [80, 81, 82, 83, 84]
@@
Figure 6: PBCs, the trajectory averaged density distribution evolution with respect to rescaled time for $L=128$ and the time step is $0.02$ (a) previous feedback, $\Delta=0$; (b) edge feedback, $\Delta=0$; (c) previous feedback, $\Delta=0.6$; (d) edge feedback, $\Delta=0.6$.
@@
Figure 7: PBCs, the trajectory averaged bipartite entanglement entropy $\overline{S}_{L/2}$ evolves with respect to rescaled time, for different feedback and $\Delta$.
@@
Appendix B Feedback protocol with boundary effect for the tilted potential
In this section, we focus on a different feedback protocol, considering the boundary effects with PBCs. This is because a tilted chain exhibits peculiar phenomena under PBCs, behaving similarly to OBCs and showing anomalous bulk dynamics. We incorporate the edge, considering the edge feedback protocol, and the quantum jump operator is written as,
@@ Appendix B Feedback protocol with boundary effect for the tilted potential
From FIG. 6 (a) and (b), $\overline{\left\langle\hat{n}_{l}\right\rangle}$ for previous and edge feedback are the same without the tilted potential. The bipartite entanglement entropy $\overline{S}_{L/2}$ are the same for different feedback likewise, from green solid lines in FIG. 7. Comparing FIG. 6 (c) and (d), despite some similarities, the steady state in (d) is not completely occupied in the left part, and the right part shows a sparse density distribution. It is evident from the orange solid lines in FIG. 7 that the $\overline{S}_{L/2}$ for the edge feedback is consistently higher than the other. The reason for the phenomenon is that edge feedback affects the pseudo-edge effect, which is caused by the tilted potential. A minority of particles can cross the pseudo edge from the first to the last site with the aid of the quantum jump with the edge feedback. If $\Delta=0$, the absence of the pseudo edge occurs, so observables remain the same for different feedback.

### v2

@@ A.1 Numerical simulation for evolution of the state
Now we consider the continuous measurement with feedback and numerical simulation of the quantum jump, i.e.,
$\left|\psi\left(t\right)\right\rangle\longrightarrow\frac{\hat{L}_{l}\left|\psi\left(t\right)\right\rangle}{\sqrt{\langle\hat{L}_{l}^{\dagger}\hat{L}_{l}\rangle}}.$ (8)
The quantum jump probability can be written as
$p_{l}=\langle\hat{L}_{l}^{\dagger}\hat{L}_{l}\rangle\gamma\delta t.$ (9)
Using Wick’s theorem,
$\langle\hat{d}_{l}^{\dagger}\hat{d}_{l}\hat{d}_{l}^{\dagger}\hat{d}_{l}\rangle=\langle\hat{d}_{l}^{\dagger}\hat{d}_{l}\rangle\langle\hat{d}_{l}^{\dagger}\hat{d}_{l}\rangle+\langle\hat{d}_{l}^{\dagger}\hat{d}_{l}\rangle\langle\hat{d}_{l}\hat{d}_{l}^{\dagger}\rangle,$ (10)
and the fermionic anticommutation relations,
$\{\hat{c}_{i},\hat{c}_{j}^{\dagger}\}=\delta_{ij},$ (11)
we get the probability of quantum jump,
$p_{l}=\langle\hat{d}_{l}^{\dagger}\hat{d}_{l}\rangle\gamma\delta t=\left\langle\psi(t)\right|\hat{d}_{l}^{\dagger}\hat{d}_{l}\left|\psi(t)\right\rangle\gamma\delta t=\|\hat{d}_{l}\left|\psi(t)\right\rangle\|^{2}\gamma\delta t.$ (12)
Figure 6: PBCs, the trajectory-averaged density distribution evolution with respect to rescaled time for $L=128$ and the time step is $0.02$ (a) previous feedback, $\Delta=0$; (b) edge feedback, $\Delta=0$; (c) previous feedback, $\Delta=0.6$; (d) edge feedback, $\Delta=0.6$. In panel (e), $\overline{S}_{L/2}$ evolves with respect to rescaled time for different feedback and $\Delta$.
Note that the linear expansion of $\hat{d}_{l}$ can be written as
$\hat{d}_{l}=\sum_{l}^{L}{[\textbf{d}]_{l}\hat{c}_{l}},$ (13)
where d is a row vector derived from the relation
$\hat{d}_{l}=\frac{\hat{c}_{l}+i\hat{c}_{l+1}}{\sqrt{2}}.$ (14)
To simulate the quantum jump, we can find a single-particle state which satisfies $\textbf{d}U_{n}\neq 0$ and $n>1$, then move $U_{n}$ to the first due to gauge freedom and transform the rest columns of $U(t)$ accordingly,
$U_{n}^{\prime}=U_{n}-\frac{\textbf{d}U_{n}}{\textbf{d}U_{1}}U_{1}.$ (15)
@@
$U^{\prime}\left(t\right)=\begin{bmatrix}{\textbf{d}}^{\dagger}&U_{2}^{\prime}&U_{3}^{\prime}&\cdots&U_{N}^{\prime}\end{bmatrix}.$ (16)
Finally, we use the QR decomposition for $e^{i\pi\mathcal{N}_{l+1}}U^{\prime}(t)$, and $\mathcal{N}_{l+1}$ is the matrix form of particle number operator.
A.2 trajectory-averaged observables
@@
When calculating the bipartite entanglement entropy $S_{L/2}$, we choose the sub system $[1,L/2]$ and then obtain all the eigenvalues $\lambda_{i}$} of sub-correlation matrix $C_{ij}^{L/2}$. $S_{L/2}$ is given by  [89, 90, 91, 92, 93]
@@
Figure 7: $\overline{S}_{L/2}$ versus rescaled time $t/\tau$. (a) $\Delta$ = $1.6$; (b) $\Delta$ = $1.8$; (c) $\Delta$ = $2.8$; (d) $\Delta$ = $3.0$. The insets show the $\overline{S}_{L/2}$ versus $\log{L}$, either at different time points or averaged over time intervals.
@@
When calculating the mutual information $I_{AB}$, we first define two subsystems, each of size $L/8$, separated by a distance of $L/2$, and label them $A$ and $B$. Using the same method previously employed for computing $\overline{S}_{L/2}$, we then determine their individual entanglement entropies, $S_{A}$ and $S_{B}$, as well as the entanglement entropy of the combined subsystem, $S_{AB}$. The mutual information is given by $I_{AB}=S_{A}+S_{B}-S_{AB}$.
When calculating the $f_{\text{skin}}$, we use the matrix $U(t)$ , which describes $|\psi(t)\rangle$, and the matrix $U_{\text{skin}}$ , which describes $|\psi_{\text{skin}}\rangle$, to perform the following calculation: $|U_{\text{skin}}^{\dagger}U(t)|^{2}$.
When computing $\|D\|_{F}$, we select the submatrix $D=C$[$L/2$: $L$, $1$: $L/2$] from the correlation matrix , and then calculate the Frobenius norm of $D$.
Figure 8: $\overline{S}_{L/2}$ evolves under different $\theta$. The green solid line represents $\Delta$ = 00, and the orange solid line represents $\Delta$ = $0.6$. The color gradient from light to dark corresponds to increasing $L$ values from smaller to larger. (a) $\theta=0.6\pi$; (b) $\theta=0.8\pi$. The inset in panel (a) provides an enlarged view of the region surrounding the transition point at $\Delta=0$.
Appendix B Feedback protocol with boundary effect for the Wannier-Stark ladder
In this section, we focus on a different feedback protocol, considering the boundary effects with PBCs. This is because a chain with Wannner-Stark ladder exhibits peculiar phenomena under PBCs, behaving similarly to OBCs and showing anomalous bulk dynamics. We incorporate the edge, considering the edge feedback protocol, and the quantum jump operator is written as,
@@ Appendix B Feedback protocol with boundary effect for the Wannier-Stark ladder
From FIG. 6 (a) and (b), $\overline{\left\langle\hat{n}_{l}\right\rangle}$ for previous and edge feedback are the same without the Wannier-Stark ladder. The bipartite entanglement entropy $\overline{S}_{L/2}$ are the same for different feedback likewise, from green solid lines in FIG. 6 (e). Comparing FIG. 6 (c) and (d), despite some similarities, the steady state in (d) is not completely occupied in the left part, and the right part shows a sparse density distribution. It is evident from the orange solid lines in FIG. 6 (e) that the $\overline{S}_{L/2}$ for the edge feedback is consistently higher than the other. The reason for the phenomenon is that edge feedback affects the pseudo-edge effect, which is caused by the Wannier-Stark ladder. A minority of particles can cross the pseudo edge from the first to the last site with the aid of the quantum jump with the edge feedback. If $\Delta=0$, the absence of the pseudo edge occurs, so observables remain the same for different feedback.
Appendix C The edge of dynamical phases in phase diagram
In this section, we discuss the ambiguous boundary between the log-law phase and the upper area-law phase ($\Delta$ = $1.6$, $1.8$), as well as the top edge of the phase diagram ($\Delta$ = $2.8$, $3.0$). In both FIG. 7 (a) and (b), the transition from the log-law phase to the area-law phase occurs at significantly later times compared to the cases of $\Delta$ = $0.0$ and $0.6$, seeing FIG. 2 (e). The insets detail the behavior before and after the crossing point, clearly illustrating the transition. In FIG. 7 (c) and (d), no transition is observed within the extended time range we selected. The insets present data averaged over various time intervals, indicating that $\overline{S}_{L/2}$ essentially remains unchanged as $\log{L}$ increases.
Subsequently, we employ the fitting function $\overline{S}_{L/2}=a\log{L}+b$ to analyze the $\overline{S}_{L/2}$ versus $\log{L}$ data for various $\Delta$ values with the aim of determining the coefficient $a$. For $\Delta=0.0$, $0.2$, $0.4$, $0.6$, $1.6$, and $1.8$, when $t/\tau$ is set to $0.5$, $0.5$, $0.5$, $0.5$, $3.1$, and $3.6$, respectively, we obtain $a\approx 0.33$, $0.43$, $0.42$, $0.37$, $0.16$, and $0.13$. However, when applying the same fitting procedure to the data for $\Delta=2.8$ and $3.0$, the resulting $a$ values are all less than $10^{-2}$, and the plot of $\overline{S}_{L/2}$ versus $\log{L}$ shows only minimal variation. We note that this behavior is similar to that observed for $\Delta=0.0$ at $t/\tau=1.8$, where the system is in the area-law phase. Therefore, combined with the previously noted observation that no transition was observed over an extended time period, we suggest that for large $\Delta$, the system remains in the area-law phase.
Appendix D The effect of variations in $\theta$ on entanglement relaxation.
In this section, we attempt to use feedback with different $\theta$ values under OBCs to investigate the effect of $\theta$ variations on the evolution of $\overline{S}_{L/2}$. In FIG. 8 (a), with $\theta=0.6\pi$, when $\Delta=0$ the corresponding $t_{c}/\tau$approx $0.98$; however, when $\Delta=0.6$, based on the continuously decreasing $\overline{S}_{L/2}$ over time, it can be inferred that its transition point will occur after $t_{c}/\tau$ $=$ $10$. In FIG. 8 (b), with $\theta=0.8$, when $\Delta=0$, $t_{c}/\tau\approx 0.83$; however, when $\Delta=0.6$, the time corresponding to $t_{c}/\tau$ is significantly larger than that of the orange lines in FIG. 2 (e). This indicates that the system in the Wannier-Stark ladder is more sensitive to variations in $\theta$.